# Chiral nature of current flow in the regime of the quantized Hall effect

### v2

@@ Abstract
The integer quantum Hall effect (QHE) belongs to the most fundamental phenomena of solid state physics and has an important application as resistance standard. It serves as a basis to understand the fractional, anomalous or spin QHEs, candidates for applications in quantum technology due to their topological properties. For optimizing all these applications it is essential to understand the current distribution inside the Hall bar, a question disputed for decades. We perform multiterminal current measurements on a Hall bar and compare the results with limiting models. We show, based on these simple experiments, that the current flow is chiral for the plateaus of quantized Hall resistance. Everywhere else, the current is described by the Drude model.
@@ I Introduction
Magnetic field is an axial vector and, therefore, breaks time-reversal symmetry, which, however, can be restored by the simultaneous reversal of magnetic field and time, $\vec{B}\to-\vec{B}$ and $t\to-t$. This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$. In this article, we explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the quantized Hall effect.
@@
In the classical equation of motion, $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$.
The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf. Fig. 1(a). If we define the chirality of a trajectory in analogy to the handedness of a screw thread, the above trajectory of a ballistic electron is left-handed (and would be right handed for a positive charge).
@@
Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively. (a) Ballistic limit (free electron). The spiral is the classical trajectory consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity. (b) Diffusive limit. For $|\mu B|<1$ the spiral motion is damped. However, the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ and tilts it by the Hall angle $\varphi$. (c) Diffusive Hall bar in the steady state limit with edges parallel to the applied $\vec{E}$ (red arrow). The Hall voltage compensates the Lorentz force, $-e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$ (or $E_{y}=|\vec{F}_{\text{L}}|/e$), such that the drift velocity is $\vec{\upsilon}_{\text{d}}$ as for $\vec{B}=0$.
@@
The Drude model describes magnetotransport in the classical limit of the diffusive regime. Its key parameter is the electron mobility $\mu$. For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering. Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$. Applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$ with the momentum scattering time $\tau=m\mu/e$, the solution of the equation of motion introduced above in the diffusive limit becomes
$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,\text{ for }\vec{B}\perp\vec{E}\,.$ (1)
@@
[For $\vec{B}\not\perp\vec{E}$ there would be a third summand inside the parantheses, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$.] For finite $B$, $\vec{j}$ is bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$ in respect to $\vec{E}$, cf. Fig. 1(b). Importantly, as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous, too.
@@
While the Drude model is designed to describe the limit of $\mu B\lesssim 1$, its prediction for the drift velocity for $\mu B\to\infty$ where the friction forces in Eq. (1) become negligible, namely $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, is identical to the ballistic solution. Since for $\mu B\to\infty$ the rapid cyclotron motion averges out over reasonable length scales, both models predict identical current densities, too, for $\mu B\to\infty$.
@@
In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$. This is a common scenario realized in many mesoscopic magnetotransport measurements.
@@
To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf. Fig. 1(c). In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons. Applied to Eq. (1) $j_{y}=0$ yields $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$. Physically, the transversal field $E_{y}$ is sourced by charge accumulated at the sample edges. With $E_{y}=\vec{\nabla}V_{\text{H}}$ it also gives rise to the Hall voltage $V_{\text{H}}$, that can be measured across the Hall bar. Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $-eE_{y}$ (the forces associated with $B$ and $E_{y}$ cancel each other), inside the Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied. Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is uniform and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
@@
While the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field and the breaking of the time-reversal symmetry are manifest in the edge charges and the related Hall voltage. In addition, the Hall resistance is non-local as it does not cause any dissipation inside the Hall bar: At $B=0$ the two terminal current measurement of a Hall bar is the sum of the ohmic sample resistance and the resistance of the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$. At finite $B$, the non-local Hall resistance enhances the resistance of the Hall bar, such that $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
@@
As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect (QHE). Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}=0$). The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance, in agreement with the classical Drude prediction for $\mu B\to\infty$.
@@
To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of chiral, one-dimensional (1D) edge channels, such that backward scattering is suppressed [1, 2]. The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$. The filling factor $\nu$ describes the number of edge channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk. The LBP assumes that the edge channels form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2]. To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization. The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential. This yields chiral current flow, more precisely, current flow in opposite directions along opposite edges.
@@
The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE. While its merit is its simplicity, its predictions are limited due to the phenomenological character as well as its neglect of interactions and non-equilibrium effects. In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4]. Hence, neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels. This local polarization corresponds to an unrealistically large Coulomb energy. Another weak point of the LBP is the coexistence of occupied and unoccupied states inside the current carrying edge channels at the chemical potential. Even, if we assume chiral edge channels, an almost complete suppression of carrier scattering, as indicated by the accuracy of the measurements of the von Klitzing constant, is unlikely under any realistic conditions.
@@
The screening theory [3, 4, 5, 6, 7, 8, 9] offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$. By taking into account the direct Coulomb interaction between electrons, it overcomes the limitations of the LBP. The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided. Early purely classical formulations considered this screening based on the Thomas-Fermi approximation [3, 4]. They demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs). The term ICS denotes areas of the 2DES, where the chemical potential lies within a local energy gap between completely filled and empty Landau levels. For this reason, at low enough temperature, inside the ICSs carriers cannot scatter and screening of the electric field is absent. ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is aligned with the chemical potential. Hence, inside compressible regions, carriers can scatter. Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions. The early calculations already show, that both, width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$. Later, for given boundary conditions the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [7, 8, 9] (later confirmed in a full self-consistent approach of the quantum- electrostatic problem [10]). The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region centered in the Hall bar [7, 8, 9]. These predictions clearly confirmed the electrostatic landscape measured in scanning probe spectroscopy experiments as a function of the magnetic field [11, 12, 13].
@@
According to the screening theory the electric field vanishes inside the perfectly screened compressible regions while it is exclusively contained inside the ICSs. For suppressed scattering the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$. Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [4, 14, 9, 15]. In case of a bulk ICS, current should then flow in the center of the Hall bar. Recently, we confirmed the predicted bulk current directly by means of an additional tiny ohmic contact in the center of a Hall bar [16]. Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [9] and confirmed in measurements [11, 12, 13, 16] are in direct contrast to the assumptions of the LBP. Therefore, it is important to find out, whether another assumption of the LBP, which is often used, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance.
@@
Even within the screening theory this is still a controversal question [3, 4, 9, 10, 15]. In the present article, we explore the question of chirality experimentally by employing multiterminal current measurements. Thereby, we probe the division of the current into various ohmic contacts of the Hall bar. Clearly, we do not directly probe the local current density distribution. Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it. To quantify this dependence, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero. In comparison, non-chiral current flow requires additional non-zero transmission coefficients, which would ultimately yield a different divison of the current into the various contacts.
Here, we show that the measured current division can be described for arbitrary magnetic fields by either one of two limiting models, introduced in Sec. II below. In Sec. III, we then present our three- and four-terminal current experiments, compare them with the predictions of the two limiting models and discuss our results.
@@
Our first model generically assumes coherent and chiral current flow. The condition of coherence is technical. In practice, it is sufficient to assume the absence of scattering. The second model applies the classical Drude model to a multiterminal current measurement. It describes diffusive and non-chiral current.
Our experiments in Sec. III show, that the generic chiral model applies for the magnetic field regions of the plateaus of quantized Hall resistance, while the Drude model describes the situation for all other magnetic fields.

### v3

# Nature of current flow in the regime of the quantum Hall effect

@@ Abstract
The integer quantum Hall effect (QHE) belongs to the most fundamental phenomena of solid state physics and has an important application as resistance standard. It serves as a basis to understand the fractional, anomalous or spin QHEs, candidates for applications in quantum technology due to their topological properties. For optimizing all these applications it is essential to understand where the current flows inside the Hall bar, a question disputed for decades. We perform multiterminal current measurements on a Hall bar and compare the results with limiting models. We confirm, based on these experiments, that the current flow is chiral for the plateaus of quantized Hall resistance. Everywhere else, including the ranges between plateaus, our results comply with the Drude model, which predicts homogeneous current flow across a homogeneous Hall-bar.
@@ I Introduction
Magnetic field is an axial vector and, therefore, breaks time-reversal symmetry, which, however, can be restored by the simultaneous reversal of magnetic field and time, $\vec{B}\to-\vec{B}$ and $t\to-t$. This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$. In the regime of the quantum Hall effect (QHE), the situation is more complicated because of an inhomogeneous distribution of the electric field in case of a quantized Hall resistance. As a result, the current density distribution in the quantum Hall regime has been a matter of discussion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] since the discovery of the QHE [21].
In this article, we use multiterminal measurements to explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the QHE. Multiterminal measurements are not capable of directly detecting the current density distribution, but based on a careful analysis they can nevertheless reveal the nature of current flow, e.g., if it is chiral or not. Before we discuss our experimental results, we introduce relevant theoretical aspects and models for comparison with our measurements.
@@
In the classical equation of motion (EOM), $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge e-e and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to a magnetic field and an electric field $\vec{E}$.
[…]
@@
Figure 1: Motion (expressed by drift velocity) of an electron confined to the xyxy-plane with in-plane electric and perpendicular magnetic field, $\vec{E}\perp\vec{B}$ (red arrows). (a) Ballistic limit (free electron): The spiral indicates the cyclotron motion of an electron consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}$ and a (left-handed) circular motion with its radius being proportional to the initial velocity. (b) Diffusive limit: the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ compared to its absolute value for $B=0$ and tilts it by the Hall angle $\varphi$. The circular motion of the free electron is damped. (c) Diffusive Hall bar with edges parallel to the applied $\vec{E}$: The Hall voltage compensates the Lorentz force, $e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$, the drift velocity $\vec{\upsilon}_{\text{d}}$ is parallel to the sample edges.
@@
Next, we consider the diffusive regime, which is dominated by momentum scattering. In its classical limit, its electron dynamics can be described by the Drude model with the electron mobility $\mu$ as central parameter. For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), momentum scattering counteracts the circular motion. Consequently, the Drude model omits the circular motion and replaces the velocity in the EOM (above) by the uniform drift velocity $\dot{}\vec{\upsilon}\to\dot{}\vec{\upsilon}_{\text{d}}$. After applying the relaxation time ansatz $\dot{\vec{\upsilon}}_{\text{d}}=-\vec{\upsilon}_{\text{d}}/\tau$, where $\tau=m\mu/e$ is the momentum scattering time, the EOM becomes a linear vector equation. Its solution,
$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left[\vec{E}-\vec{E}\times\mu\vec{B}+\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}\right]\,$, (1)
@@
can be simply determined by rearranging the EOM. For $\vec{B}\perp\vec{E}$ the term proportional to $\vec{E}\cdot\vec{B}$ vanishes and $\vec{\upsilon}_{\text{d}}$ is confined to a plane perpendicular to $\vec{B}$. Compared to $\vec{\upsilon}_{\text{d}}=-\mu\vec{E}$ for $B=0$, a finite magnetic field reduces its absolute value according to $|\vec{\upsilon}_{\text{d}}|=\mu E/\sqrt{1+(\mu B)^{2}}$. In respect to the direction of $\vec{E}$, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, is then bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$, cf. Fig. 1(b). Importantly, as long as both the carrier density, $n_{\text{s}}$, and the mobility, $\mu$, are homogeneous, $\vec{j}$ is homogeneous, too.
@@
As argued above, the Drude model applies for $\mu B\lesssim 1$. Interestingly, for the opposite limit, namely $\mu B\to\infty$ and $R_{\text{c}}\ll l_{\text{m}}$, the rapid circular motion of the cyclotron trajectory is not damped but averages out over reasonable length scales. At the same time, for $\mu B\to\infty$, the friction forces in Eq. (1) become negligible, such that the current density predicted by the Drude model becomes $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}\to-en_{\text{s}}\vec{E}\times\vec{B}/B^{2}$, which is independent of $\mu$ and identical to the solution for ballistic electrons.
@@
In the remainder of the article, we restrict ourselves to a degenerate two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$. This is a common scenario realized in many mesoscopic magnetotransport measurements.
@@
To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf. Fig. 1(c). In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons. Applying $j_{y}=0$ to Eq. (1) yields the classical Hall effect with $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$, where we introduced the Hall resistance $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$. The transversal field $E_{y}$ is generated by charge accumulated at the sample edges, which can be measured in terms of the Hall voltage $V_{\text{H}}$, related to $E_{y}$ via $E_{y}=|\vec{\nabla}V_{\text{H}}|$. Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $eE_{y}$, the Drude model predicts that inside a Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied. Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is then predicted to be uniform and the current becomes $I=j_{x}b$, with $b$ being the width of the Hall bar.
@@
Although the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field is actually manifest in the edge charges and the related Hall voltage, which break the time-reversal symmetry. Interestingly, other than an ohmic resistance, the Hall resistance does not add dissipation inside the Hall bar. (Remember, the related transversal field compensates the Lorentz force.) Instead, the Hall resistance is non-local. It causes additional dissipation where the current flow becomes parallel to $\vec{\nabla}V_{\text{H}}$, which typically happens near or inside the current carrying contacts. (For the quantum Hall regime, an enhanced dissipation near the contacts was experimentally confirmed [22, 23]). For a constant current, the overall dissipation is proportional to the two-terminal resistance, $R_{\text{2T}}$, of the Hall bar sample. At $B=0$, it is the sum of the resistances of the Hall bar and the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$. At finite $B$, the resistance can be described as $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
@@
As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect. Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}\to 0$). The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance. The related Shubnikov-de-Haas (SdH) oscillations of the longitudinal resistance reveal in the regime of the QHE an oscillating $R_{0}(B)$ with growing maxima between the quantized plateaus as $B$ is increased, cf. Fig. 13 in Appendix C.
@@
To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of chiral, one-dimensional (1D) edge channels, such that backward scattering is suppressed [2, 3]. The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$ [24], which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$. The filling factor $\nu$ describes the number of edge channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk. The LBP assumes that the edge channels form where the Landau levels, whose potential energies increase towards the depleted sample edges, intersect with the Fermi level [2, 3]. To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to localization in a disorder potential in a high magnetic field [3]. The direction of the current flow in the 1D edge channels is predetermined by the gradient of the confinement potential. This yields current flow in opposite directions along opposite edges. Following the usual practice, we call this chiral current flow. Depending on the direction of the magnetic field the chiral current flow is clockwise or counter-clockwise.
@@
The LBP provides an intuitive explanation for the quantization of the Hall resistance and the accompanying vanishing longitudinal resistance of the QHE. While its merit is its simplicity, it neglects interactions and non-equilibrium effects, which gives it a phenomenological character with a limited range of validity. For example, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [5]. It means, that neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels. This local polarization would give rise to an unrealistically large Coulomb energy. Another problem is the coexistence of occupied and unoccupied states at the chemical potential inside the current carrying edge channels according to the LBP. Even, if we take into account chiral current flow, an almost complete suppression of carrier scattering, as indicated by the high accuracy of the measurements of the von Klitzing constant [25], would then be unlikely under realistic conditions.
@@
The screening theory [5, 26, 27, 28, 29, 30, 14] offers an alternative explanation of the QHE. By taking into account the direct Coulomb interaction between electrons, it overcomes the limitations of the LBP. It is based on the realization, that the Coulomb interaction of free carriers allows screening of the local electric field by means of rearrangement of carriers. This screening helps to avoid the highly charged stripes and thereby lower the free energy. However, the screening is hindered, if Landau-quantization in a strong magnetic field causes an energy gap. Early purely classical formulations included screening based on the Thomas-Fermi approximation [5, 26]. By minimizing the free energy, they demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs). The term ICS denotes areas of the 2DES, where the chemical potential lies within a local energy gap between completely filled and empty Landau levels. For this reason, at low enough temperature, inside the ICSs carriers cannot scatter and screening of the electric field is absent. ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is exactly aligned with the chemical potential. Hence, inside compressible regions, carriers can scatter. Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions [5]. The early calculations already show, that both, width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$. Later, for given boundary conditions, the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [29, 30, 14] (later confirmed in a fully self-consistent approach of the quantum-electrostatic problem [31]). The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region (bulk ICS) centered in the Hall bar [29, 30, 14]. These calculations confirmed the electrostatic landscape measured in high-resolution scanning probe spectroscopy experiments as a function of the magnetic field [12, 32, 33], which where indicated earlier by scanning electro-optical (Pockels) effect experiments in larger samples [4, 34].
@@
According to the screening theory the electric field vanishes inside the perfectly screened compressible regions while it is exclusively contained inside the ICSs. For suppressed scattering ($\mu B\gg 1$) the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ (ballistic electrons). Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [26, 35, 14]. In case of a bulk ICS, current should then flow near the center of the Hall bar, while at the same time the Hall resistance should remain quantized. Recently, we precisely confirmed this predicted bulk current by exploring current flow through an additional tiny ohmic contact in the center of a Hall bar [36]. Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [14] and confirmed in measurements [4, 34, 12, 32, 33, 36] are in direct contrast to the assumptions of the LBP. Therefore, it is important to find out, whether another assumption of the LBP, based on the axial symmetry of the magnetic field vector, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance and whether the chirality survives even in the case of bulk current.
The distribution of the local current density in the quantum Hall regime has been controversally discussed for some time, in particular, in articles going beyond the LBP [5, 26, 9, 8, 35, 14, 17, 20, 31, 36]. In the present article, we present experiments designed to explore the question of the chirality of the current flow by employing multiterminal current measurements. In detail, we probe the division of the current into various ohmic contacts of the Hall bar. Measuring currents into contacts, we cannot directly probe the local current density distribution. Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it, see Sec. III.3 for related measurements.
@@
To quantify the current distribution into various contacts, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero. In comparison, non-chiral current flow requires additional non-zero transmission coefficients. Clearly, the two cases lead to different distributions of the current into the various contacts.
In Sec. II we introduce the two limiting models, the first one describing coherent and chiral current flow and the second one describing diffusive current (Drude model for multiple terminals). In Sec. III, we then present our three- and four-terminal current experiments and show that the measured current distribution into various contacts can be described for arbitrary magnetic fields by either one of two limiting models. Our results suggest, that the current flow is chiral for the quantized plateaus but unidirectional between the plateaus.
@@
Our first model generically assumes coherent and chiral current flow. The condition of coherence is assumed for convenience. While this assumption is technically not necessary, the absence of scattering in an ICS would actually imply coherent current flow. The second model applies the classical Drude model to a multiterminal current measurement. It describes incoherent, diffusive and non-chiral current.